# EgoVideo: Exploring Egocentric Foundation Model and Downstream Adaptation

### v1

@@ 1 Introduction
To tackle the challenge, we propose a 3-stage training paradigm for egocentric video understanding, including action recognition, domain adaptation, and multi-instance retrieval. Specifically, we first adopt a video foundation model [31] that is pre-trained on large-scale video-language datasets [30]. With the help of rich vision features and wide range of action-aware knowledge, this model is capable of extracting general video feature representations, acting as a good start point for subsequent feature learning. In the second stage, to mitigate the domain gap between web-scale video datasets and egocentric videos, we perform post-training on a collection of egocentric videos [12, 23, 27], effectively transferring the general video feature representations to egocentric domain. We term this model as EgoVideo, consisting of a strong egocentric video encoder EgoVideo-V and a text encoder EgoVideo-T. In the third stage, we conduct task-specific fine-tuning of EgoVideo-V and EgoVideo-T on three different action understanding tasks, i.e., action recognition, domain adaptation action recognition, and multi-instance retrieval.
Experimental results show that our 3-stage strategy has led to a remarkable improvement in overall model performance. The model excels at understanding fine-grained, action-specific information, demonstrating strong performance on action recognition and multi-instance retrieval. Moreover, benefiting from the multi-stage training, our model exhibits video understanding ability across wide range of actions.
@@
2 Training Process of EgoVideo
2.1 Stage1: General Video Pre-training
2.2 Stage2: Egocentric Video Post-training
To transfer the video foundation model learned in the general video domain into the egocentric domain, we initialize the model with weights pre-trained in stage 1 and conduct post-training. Specifically, we collect a broad range of paired egocentric video-text pairs from public video datasets, such as Ego4d [12], HowTo100M [23], EgoExoLearn [17], and Ego4d GoalStep [27]. This results in around 7M video-text pairs. We train the model for 5 epochs on the hybrid data to improve the egocentric video understanding ability. The model is optimized via a standard visual-text contrastive loss. During training, we also examine the model’s egocentric video understanding ability on EPIC-Kitchen-100 zero-shot multi-instance retrieval benchmark [5], and the results are shown in Table 3.6. We term this egocentric video foundation model as EgoVideo, consisting of a strong egocentric video encoder EgoVideo-V and a text encoder EgoVideo-T.
2.3 Stage 3: Egocentric Downstream Adaptation
After stage 2 training, we obtain a video foundation model EgoVideo tailored for egocentric domain. We use this model to initialize the models in stage 3. In this stage, we conduct task-specific fine-tuning on the training sets. We put the detailed task-specific fine-tuning process of each task in the following section.
3 Task-specific Finetuning
3.1 Task 1: Natural Language Queries @ Ego4D
@@
3.2 Task 2: GoalStep - Step Grounding @ Ego4D
[…]
3.3 Task 3: Moment Queries @ Ego4D

### v2

@@ 1 Introduction
Figure 1: The workflow of the training process of EgoVideo. It includes 3 stages: in the first stage, we filter and select high-quality egocentric video and text pairs from multiple existing datasets. Then we perform post pertaining using the data in stage 1 by standard video-text contrastive learning. Finally, we adapt the pretrained EgoVideo model to different downstream tasks.
To tackle the challenge, we propose a 3-stage training paradigm for egocentric video understanding, including multiple tasks like natural language grounding, domain adaptation, and multi-instance retrieval. Specifically, we first filter and select high-quality egocentric video and text pairs from multiple existing datasets [12, 23, 27, 17]. These high-quality data serve as the foundational data for transferring models learned from general domains to the egocentric domain. We adopt a video foundation model [31] that is pre-trained on large-scale video-language datasets [30]. With the help of rich vision features and a wide range of action-aware knowledge, this model is capable of extracting general video feature representations, acting as a good starting point for subsequent feature learning. In the second stage, to mitigate the domain gap between web-scale video datasets and egocentric videos, we perform post-training on the selected data, effectively transferring the general video feature representations to egocentric domain. We term the resulting model as EgoVideo, consisting of a strong egocentric video encoder EgoVideo-V and a text encoder EgoVideo-T. In the third stage, we conduct task-specific fine-tuning of EgoVideo-V and EgoVideo-T on three different egocentric video understanding tasks, e.g., natural language queries, domain adaptation action recognition, and multi-instance retrieval.
Experimental results show that our 3-stage strategy has led to a remarkable improvement in overall model performance. The model excels at understanding fine-grained, action-specific information, demonstrating strong performance in action recognition and multi-instance retrieval. Moreover, benefiting from the multi-stage training, our model exhibits video understanding ability across a wide range of actions.
2 Training Process of EgoVideo
2.1 Stage1: Augmented Data Selection
To better transfer the video foundation model learned in the general video domain into the egocentric domain, we collect a broad range of paired egocentric video-text pairs from public video datasets, such as Ego4d [12], HowTo100M [23], EgoExoLearn [17], and Ego4d GoalStep [27] by automatic filtering techniques. We do this to ensure a wider range of egocentric data and maintain the pertaining data quality. This results in around 7M video-text pairs.
2.2 Stage2: Egocentric Video Post-training
We then perform the post-pretraining process and train the model for 5 epochs on the hybrid data in Stage 1 to improve the egocentric video understanding ability. The model is optimized via a standard visual-text contrastive loss. During training, we also examine the model’s egocentric video understanding ability on EPIC-Kitchen-100 zero-shot multi-instance retrieval benchmark [5], and the results are shown in Table 3.6. We term this egocentric video foundation model as EgoVideo, consisting of a strong egocentric video encoder EgoVideo-V and a text encoder EgoVideo-T.
2.3 Stage 3: Egocentric Downstream Adaptation
After stage 2 training, we obtain a video foundation model EgoVideo tailored for the egocentric domain. We use this model to initialize the models in stage 3. In this stage, we conduct task-specific fine-tuning on the training sets. We put the detailed task-specific fine-tuning process of each task in the following section.
3 Task-specific Finetuning
3.1 Task 1: Natural Language Queries @ Ego4D
@@
3.2 Task 2: GoalStep - Step Grounding @ Ego4D
[…]
3.3 Task 3: Moment Queries @ Ego4D